# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

### v1

@@ III-B Path planing based on tangent cones
According to Nagumo’s theorem, it is sufficient to address the collision-free path planning problem by solving the following constrained optimization problem
@@
$\begin{split}\min\limits_{\bm{\tau}}&~{}\|\bm{\tau}-\bm{\kappa}_{0}(\mathbf{x})\|,\\ \text{s.t.}&~{}\bm{\tau}\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},\end{split}$ (12)
@@
where $\bm{\kappa}_{0}(\mathbf{x})$ is a nominal control law for motion-to-goal and is designed here as the following smooth function:
$\bm{\kappa}_{0}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (13)
with $k_{0}(\mathbf{x})>0$ given by
$k_{0}(\mathbf{x}):=\dfrac{\alpha}{\sqrt{\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}+\beta^{2}}},$ (14)
where $\alpha$ and $\beta$ are positive design constants. From (13) and (14), it follows that $\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, which allows us to generate a saturated nominal vector field. The problem (12) is clearly a nearest point problem, which has an explicit solution that will be demonstrated in the following.
Figure 4: The resulting smooth navigation trajectories maintains a safety margin of size $\epsilon$ (light blue regions) from the workspace boundary and obstacles. The dashed circles denote the influence regions from each obstacle. Moreover, the orange points are the unstable stationary points associated with each obstacles.
@@
The robot workspace $\mathcal{W}^{\ast}$ is a convex set, so does $\mathcal{W}^{\epsilon}$, which together with the fact that $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$ suggests that for all $\mathbf{x}\in\partial\mathcal{W}^{\epsilon}$, $\bm{\kappa}_{0}(\mathbf{x})$ points inside the free space $\mathcal{X}_{\epsilon}$ (i.e., $\bm{\kappa}_{0}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x})$) and thus is a solution to (12). Moreover, for all $\mathbf{x}\in\text{int}(\mathcal{X}_{\epsilon})$, it follows that $\bm{\kappa}_{0}(\mathbf{x})$ is also the solution to (12). Next we check the obstacle boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$. As $\partial\mathcal{O}^{\epsilon}$ is a smooth hypersurface of $\mathbb{R}^{2}$, which is orientable, there exists a continuously differentiable map (known as the Gauss map [28]) $\mathbf{n}:\partial\mathcal{O}^{\epsilon}\to\mathbb{S}^{1}$ such that for all $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$, $\mathbf{n}(\mathbf{x})$ is the outward unit normal vector to $\partial\mathcal{O}^{\epsilon}$. As clearly seen in Fig. 2, the tangent cone at any $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ is a half-space
$\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})\leq 0\},$ (15)
which is a convex function. Thus, (12) has a unique solution. Based on the above argument, it is known that if $\mathbf{x}\in\mathcal{X}_{\epsilon}\setminus\partial\mathcal{O}^{\epsilon}$ or $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, then $\bm{\kappa}_{0}(\mathbf{x})$ is a solution to (12). On the other hand, if $\mathbf{x}\in\mathcal{O}^{\epsilon}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$, the closest point of (12) is obtained by the orthogonal projection $\mathbf{Q}(\mathbf{n}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{n}(\mathbf{x})\mathbf{n}^{\top}(\mathbf{x})$ onto the tangent hyperplane of $\partial\mathcal{X}_{\epsilon}$, defined by $\mathbf{H}_{\partial\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})=0\}$. Then, a general solution to (12) is as follows:
$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0,\\ &\mathbf{Q}(\mathbf{n}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0.\end{split}\right.$ (16)
By inspecting (16), we find that the resulting vector field is discontinuous at some boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ like $\bm{P}_{1}-\bm{P}_{4}$ in Fig. 2. To address this issue, a continuous version of (16) is proposed in the sequel. Following the line of [4], we specify an influence region for each obstacle (marked by the dashed line in Fig. 4), which is defined as $\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}\}$, where $\epsilon^{\ast}\in(\epsilon,h]$ and $i\in\mathbb{I}$. Obstacle avoidance is activated only when the position $\mathbf{x}$ of $\bm{P}$ enters the influence regions of obstacles. To proceed, a bearing vector is defined as
$\mathbf{b}(\mathbf{x}):=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\|\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}\|}=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\text{d}_{\mathcal{O}}(\mathbf{x})}.$ (17)
@@
for all $\mathbf{x}$ with $\epsilon\leq\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$. Since $\text{d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h)$, $\forall i,j\in\mathbb{I}$, $i\neq j$ (referring to (5) in Assumption 2) and $\epsilon^{\ast}\leq h$, there can be only one obstacle $\mathcal{O}_{i}$, $i\in\mathbb{I}$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})=\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}$. With this in mind, $\mathbf{b}(\mathbf{x})$ in (17) can be computed by $\mathbf{b}(\mathbf{x})=(\mathbf{c}_{i}-\mathbf{x})/\|\mathbf{c}_{i}-\mathbf{x}\|$. Note that when $\mathbf{x}\in\partial\mathcal{O}_{\epsilon}$, $\mathbf{b}(\mathbf{x})$ is equivalent to the Gauss map $\mathbf{n}(\mathbf{x})$. Now the control law (16) is modified as
@@
$\bm{\tau}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0,\end{split}\right.$ (18)
@@
where $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in\mathbb{R}$ is a $C^{1}$ bump function that smoothly transitions from $1$ to 00 on the interval $\text{d}_{\mathcal{O}}(\mathbf{x})\in[\epsilon,\epsilon^{\ast}]$, that is to say, its first derivatives equal to zero at the endpoints of the interval $[\epsilon,\epsilon^{\ast}]$. A simple choice of $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ is
$\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=\left\{\begin{split}&1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon,\\ &\lambda(\text{d}_{\mathcal{O}}(\mathbf{x})),&\text{if}~{}&\epsilon<\text{d}_{\mathcal{O}}(\mathbf{x})<\epsilon^{\ast},\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\geq\epsilon^{\ast},\end{split}\right.$ (20)
with
$\lambda(\text{d}_{\mathcal{O}}(\mathbf{x}))=\dfrac{1}{2}\left[1-\cos\left(\pi\dfrac{\epsilon^{\ast}-\text{d}_{\mathcal{O}}(\mathbf{x})}{\epsilon^{\ast}-\epsilon}\right)\right].$ (21)
One can verify that $\phi(\epsilon)=1$, $\phi(\epsilon^{\ast})=0$, and
$\left.\dfrac{d\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))}{d\text{d}_{\mathcal{O}}(\mathbf{x})}\right|_{\text{d}_{\mathcal{O}}(\mathbf{x})=0~{}\text{or}~{}\epsilon^{\ast}}=0.$
Since the nominal control law, $\bm{\kappa}_{0}(\mathbf{x})$, the bearing vector, $\mathbf{b}(\mathbf{x})$, and the bump function, $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$, are continuously differentiable, the resulting control law $\bm{\tau}(\mathbf{x})$ in (18) is continuous and piecewise continuously differentiable on the domain $\mathcal{X}$. It is not differentiable only at those points satisfying $\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}$ and $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})=0$. Now the main result of the path planning problem is summarized in the following theorem.
@@
[…]
@@
    The free space $\mathcal{X}_{\epsilon}$ in (7) is forward invariant.
@@
    For any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the kinematics (11) admits a unique solution in forward time, which asymptotically converges to the set $\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\mathcal{S}_{i}$ with $\mathcal{S}_{i}$ given by
    $\mathcal{S}_{i}:=\left\{\mathbf{x}\in\mathcal{X}_{\epsilon}\bigg{|}{\rm d}_{\mathcal{O}_{i}}(\mathbf{x})=\epsilon,~{}\dfrac{\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})}{\|\bm{\kappa}_{0}(\mathbf{x})\|}=1\right\}.$ (22)
@@
    The undesired stationary points $\mathbf{s}_{i}\in\mathcal{S}_{i}$, associated with obstacles $\mathcal{O}_{i}$, are locally unstable and only have a stable manifold of measure zero.
  4. 4.
@@
The proposed safe path planning scheme offers an analytical solution and thus is computationally much more efficient than the optimization-based methods. Moreover, unlike the navigation functions, the scheme can achieve almost global asymptotic convergence without requiring unknown parameter tuning, making it easy to implement in practice.
@@
A smooth and saturated nominal control law (18) is proposed, instead of the traditional proportional control law $k_{0}(\mathbf{x})=-k(\mathbf{x}-\mathbf{x}^{\ast})$ [4, 21], with the aim of generating a saturated vector field (18). This is justified by the inequality $\|\bm{\tau}(\mathbf{x})\|\leq\|\bm{\kappa}_{0}(\mathbf{x})\|\leq\alpha$, where the equality $\|\bm{\Pi}(\mathbf{x})\|=1$ has been used. In fact, different types of saturation functions can be used to design $\kappa_{0}(\mathbf{x})$ as an alternative, such as the standard $\text{sat}(\cdot)$, hyperbolic tangent function $\tanh(\cdot)$, $\text{arctan}(\cdot)$, etc. However, it is worth noting that these functions may result in the nominal control direction not aligning directly with the desired position, which in turn leads to an unnecessarily long path.
III-C Adaptive tube-following control
The planning algorithm introduced in Sec. III-B can generate a collision-free reference trajectory for the WMR. To avoid conflicts of notations, the reference trajectory is denoted as $\mathbf{x}_{d}$, which is governed by
$\dot{\mathbf{x}}_{d}=\bm{\tau}_{d},$ (23)
where $\bm{\tau}_{d}$ is given by (18) but with $\mathbf{x}$ replaced by $\mathbf{x}_{d}$. Let us define the position tracking error as $\mathbf{x}_{e}:=\mathbf{x}-\mathbf{x}_{d}$ and impose an inequality constraint on it as follows:

### v2

@@ III-B Prescribed-time path planning
A tangent cone based control policy (vector field) $\bm{\tau}=\mathbf{h}(\mathbf{x})$ is designed to achieve collision-free path planning. According to Theorem 1, we consider a nearest point problem
@@
$\begin{split}\min\limits_{\mathbf{h}}&~{}\|\mathbf{h}(\mathbf{x})-\bm{\kappa}_{0}(\mathbf{x})\|\\ \text{s.t.}&~{}\mathbf{h}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{X}_{\epsilon},\end{split}$ (8)
@@
where $\bm{\kappa}_{0}(\mathbf{x})$ is a nominal control law for motion-to-goal and is designed here as
$\bm{\kappa}_{0}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast}),$ (9)
with $k_{0}>0$ being a constant.
@@
The robot workspace $\mathcal{W}^{\ast}$ is a convex set, so does $\mathcal{W}^{\epsilon}$, which suggests that for all $\mathbf{x}\in\partial\mathcal{W}^{\epsilon}$, $\bm{\kappa}_{0}(\mathbf{x})$ points inside the free space $\mathcal{X}_{\epsilon}$ (i.e., $\bm{\kappa}_{0}(\mathbf{x})\in\mathbf{T}_{\mathcal{{X}}_{\epsilon}}(\mathbf{x})$) and thus is a solution to (8). Moreover, for all $\mathbf{x}\in\text{int}(\mathcal{X}_{\epsilon})$, $\bm{\kappa}_{0}(\mathbf{x})$ is also the solution to (8). Next we check the obstacle boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$. As $\partial\mathcal{O}^{\epsilon}$ is a smooth hypersurface of $\mathbb{R}^{2}$ and is orientable, there exists a continuously differentiable map (known as Gauss map [30]) $\mathbf{n}:\partial\mathcal{O}^{\epsilon}\to\mathbb{S}^{1}$ such that for all $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$, $\mathbf{n}(\mathbf{x})$ is the outward unit normal vector to $\partial\mathcal{O}^{\epsilon}$. As clearly seen in Fig. 2, the tangent cone at any $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ is a half-space $\mathbf{T}_{\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})\leq 0\}$, which is a convex function. Thus, (8) has a unique solution. In summary, if $\mathbf{x}\in\mathcal{X}_{\epsilon}\setminus\partial\mathcal{O}^{\epsilon}$ or $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, then $\bm{\kappa}_{0}(\mathbf{x})$ is a solution to (8). While if $\mathbf{x}\in\mathcal{O}^{\epsilon}\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$, the closest point of (8) is obtained by the orthogonal projection $\mathbf{Q}(\mathbf{n}(\mathbf{x})):=\mathbf{I}_{2}-\mathbf{n}(\mathbf{x})\mathbf{n}^{\top}(\mathbf{x})$ onto the tangent hyperplane of $\partial\mathcal{X}_{\epsilon}$, defined by $\mathbf{H}_{\partial\mathcal{X}_{\epsilon}}(\mathbf{x}):=\{\mathbf{y}\in\mathbb{R}^{2}\mid(\mathbf{y}-\mathbf{x})^{\top}\mathbf{n}(\mathbf{x})=0\}$. Then, a general solution to (8) is as follows: $\mathbf{h}(\mathbf{x})=\bm{\kappa}_{0}(\mathbf{x})$ if $\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon$ or $\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})\leq 0$, and $\mathbf{h}(\mathbf{x})=\mathbf{Q}(\mathbf{n}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$ if $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{n}(\mathbf{x})>0$. The resulting vector field is discontinuous at some boundary points $\mathbf{x}\in\partial\mathcal{O}^{\epsilon}$ like the green points in Fig. 2. To address this issue, a continuous control law is proposed in the sequel. Following the line of [4], we specify an influence region for each obstacle (marked by the dashed line in Fig. 2), which is defined as $\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}\}$, where $\epsilon^{\ast}\in(\epsilon,h]$ and $i\in\mathbb{I}$. Obstacle avoidance is activated only when the position $\mathbf{x}$ of $\bm{P}$ enters the influence regions of obstacles. To proceed, a bearing vector is defined as
$\mathbf{b}(\mathbf{x}):=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\|\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}\|}=\dfrac{\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x})-\mathbf{x}}{\text{d}_{\mathcal{O}}(\mathbf{x})},$ (10)
@@
where $\mathbf{P}_{\partial\mathcal{O}}(\mathbf{x}):=\{\mathbf{y}\in\partial\mathcal{O}\mid\|\mathbf{y}-\mathbf{x}\|=\text{d}_{\mathcal{O}}(\mathbf{x})\}$ is a set-valued map. Since $\text{d}(\mathcal{O}_{i}^{\ast},\mathcal{O}_{j}^{\ast})>2(r+h)$, $\forall i,j\in\mathbb{I}$, $i\neq j$ (referring to Assumption 2) and $\epsilon^{\ast}\leq h$, there can be only one obstacle $\mathcal{O}_{i}$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})=\text{d}_{\mathcal{O}_{i}}(\mathbf{x})\leq\epsilon^{\ast}$. With this in mind, $\mathbf{b}(\mathbf{x})$ in (10) can be computed by $\mathbf{b}(\mathbf{x})=(\mathbf{c}_{i}-\mathbf{x})/\|\mathbf{c}_{i}-\mathbf{x}\|$. Note that when $\mathbf{x}\in\partial\mathcal{O}_{\epsilon}$, $\mathbf{b}(\mathbf{x})$ is equivalent to the Gauss map $\mathbf{n}(\mathbf{x})$. Now the control law is modified as
@@
$\mathbf{h}(\mathbf{x})=\left\{\begin{split}&\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}~{}\text{or}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})\leq 0,\\ &\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x}),&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon^{\ast}~{}\text{and}\\ &&&\bm{\kappa}_{0}^{\top}(\mathbf{x})\mathbf{b}(\mathbf{x})>0,\end{split}\right.$ (11)
@@
where $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in\mathbb{R}$ is a $C^{1}$ bump function that smoothly transitions from $1$ to 00 on the interval $\text{d}_{\mathcal{O}}(\mathbf{x})\in[\epsilon,\epsilon^{\ast}]$. A simple choice of $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ is
$\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=\left\{\begin{aligned} &1,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\leq\epsilon,\\ &\lambda(\text{d}_{\mathcal{O}}(\mathbf{x})),&\text{if}~{}&\epsilon<\text{d}_{\mathcal{O}}(\mathbf{x})<\epsilon^{\ast},\\ &0,&\text{if}~{}&\text{d}_{\mathcal{O}}(\mathbf{x})\geq\epsilon^{\ast},\end{aligned}\right.$ (13)
where $\lambda(\text{d}_{\mathcal{O}}(\mathbf{x}))=0.5[1-\cos(\pi(\epsilon^{\ast}-\text{d}_{\mathcal{O}}(\mathbf{x}))/(\epsilon^{\ast}-\epsilon))]$. It is clear that the modified control law (11) is continuous and piecewise continuously differentiable on the domain $\mathcal{X}$.
@@
Consider the kinematics defined by (7). If $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$ and the obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}$ satisfying Assumption 2, then the continuous control law (11) with $\bm{\kappa}_{0}(\mathbf{x})$ given by (9) can guarantee that:
@@
    The free space $\mathcal{X}_{\epsilon}$ is forward invariant.
@@
    For any $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, the solution of (7) admits a unique solution in forward time, which asymptotically converges to the set $\mathcal{C}:=\{\mathbf{x}^{\ast}\}\cup_{i=1}^{n}\{\mathbf{s}_{i}\}$, where the stationary point $\mathbf{s}_{i}$ that satisfies ${\rm d}_{\mathcal{O}_{i}}(\mathbf{s}_{i})=\epsilon\wedge\bm{\kappa}_{0}^{\top}(\mathbf{s}_{i})\mathbf{b}(\mathbf{s}_{i})/\|\bm{\kappa}_{0}(\mathbf{s}_{i})\|=1$ is locally unstable.
@@
To further achieve prescribed-time path planning, the TSTF $t:=\eta(s)$ defined in (6) is introduced to squeeze and map the collision-free trajectory of (7) from an infinite time interval $s\in[0,\infty)$ to a prescribed finite time interval $t\in[0,T)$. Let $\eta^{\prime}(s):=d\eta(s)/ds$ and define a prescribed-time gain function (PTGF) $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ as
$\alpha(t)=\alpha(\eta(s)):=\dfrac{1}{\eta^{\prime}(s)}=\dfrac{T}{T-t},~{}t\in[0,T),$ (14)
showing that $\alpha(t)>0$ is continuously differentiable on $t\in[0,T)$ and satisfies $\alpha(0)=1$ and $\lim_{t\to T}\alpha(t)=\infty$. Here the prescribed-time control law is designed as $\bm{\tau}=\mathbf{h}^{\ast}(\mathbf{x},t)$, where $\mathbf{h}^{\ast}$ is given by
$\mathbf{h}^{\ast}(\mathbf{x},t)=\left\{\begin{split}&\alpha(t)\mathbf{h}(\mathbf{x}),&t&\in[0,T),\\ &\mathbf{h}(\mathbf{x}),&t&\in[T,\infty),\end{split}\right.$ (15)
with $\mathbf{h}(\mathbf{x})$ and $\alpha(t)$ defined in (11) and (14), respectively.
Theorem 3.
Consider the kinematics given by (7) with initial conditions satisfying $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$. Then, the closed-loop system $\dot{\mathbf{x}}=\mathbf{h}^{\ast}(\mathbf{x},t)$ converge to the set $\mathcal{C}$ at $t=T$, while avoiding obstacle regions $\mathcal{O}_{i}^{\epsilon}$, $i\in\mathbb{I}$. The system under $\bm{\tau}=\mathbf{h}^{\ast}(\mathbf{x},t)$ follows the same path as the closed-loop system $\dot{\mathbf{x}}=\mathbf{h}(\mathbf{x})$, but goes faster within the time window $[0,T)$.
Proof: The proof is conducted over two time intervals.
1) Consider the interval $t\in[0,T)$. The closed-loop system is written as
$\dot{\mathbf{x}}(t)=\alpha(t)\mathbf{h}(\mathbf{x}(t)).$ (16)
Let $\bar{\mathbf{x}}(s):=\mathbf{x}(t)=\mathbf{x}(\eta(s))$. Then, the system (16) is rewritten in the stretched infinite-time interval $s\in[0,\infty)$ as
$\bar{\mathbf{x}}^{\prime}(s):=\dfrac{d\bar{\mathbf{x}}(s)}{ds}=\dfrac{d\mathbf{x}(\eta(s))}{d\eta(s)}\cdot\dfrac{d\eta(s)}{ds}.$ (17)
In view of 14 and (16), 17 becomes
$\bar{\mathbf{x}}^{\prime}(s)=\mathbf{h}(\bar{\mathbf{x}}(s)).$ (18)
The initial conditions satisfy $\bar{\mathbf{x}}^{\prime}(0)=\dot{\mathbf{x}}(0)$ and $\bar{\mathbf{x}}(0)=\mathbf{x}(0)$, where the facts that $\eta(0)=0$ and $\alpha(0)=1$ have been used. As (18) has the same solution as the closed-loop system $\dot{\mathbf{x}}(t)=\mathbf{h}(\mathbf{x}(t))$, it follows from Theorem 2 that $\bar{\mathbf{x}}(s)$ asymptotically converges to the set $\mathcal{C}$. This implies that $\mathbf{x}(t)$ converges to the set $\mathcal{C}$ at the prescribed time $T$, due to $\mathbf{x}(t)=\bar{\mathbf{x}}(s)$ and $t\to T$ as $s\to\infty$. Additionally, Theorem 2 ensures that $\bar{\mathbf{x}}(s)$ is collision-free on $t\in[0,\infty)$. As a result, $\mathbf{x}(t)$ is also collision-free on $t\in[0,T)$.
[…]
@@
The proposed prescribed-time control policy (15) is discontinuous at $t=T$ and cannot be practically implemented due to the unboundedness of $\alpha(t)$ as $t\to T$. To remedy this, a continuous version is given as
$\tau=\mathbf{h}^{\ast}(\mathbf{x},t)=\alpha^{\ast}(t)\mathbf{h}(\mathbf{x}),$ (19)
with the PTGF given by
$\alpha^{\ast}(t):=\left\{\begin{aligned} &\alpha(t),&~{}&t\in[0,T^{\ast}),\\ &\alpha(T^{\ast}),&~{}&t\in[T^{\ast},\infty),\end{aligned}\right.$ (20)
where $T^{\ast}:=T-\varsigma>0$, and $0<\varsigma\ll T$ is a sufficiently small constant. The resulting continuous vector field can guide $\mathbf{x}$ to a small neighborhood of the set $\mathcal{C}$ at $t=T^{\ast}$. Since $\alpha^{\ast}(t)\equiv T/\varsigma>0$ on $[T^{\ast},\infty)$, $\mathbf{x}(t)$ will asymptotically converge to the set $\mathcal{C}$ after $t=T^{\ast}$. It is clear from (9) that choosing a larger $k_{0}$ will increase the convergence rate of the planner during the transient-state phase. However, this comes at the expense of increased reference velocity.
@@
Although the prescribed-time path planner (19) is partially inspired by [4], there remain substantial differences in the following two aspects:
  1. 1.
    A $C^{1}$ bump function $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))$ defined in (13) is introduced into the projection function $\mathbf{\Pi}(\mathbf{x})$, which makes the resulting vector filed more smoother;
  2. 2.
    By incorporating a PTGF $\alpha(t)$ (generated from the TSTF defined in (6)) into the planning algorithm developed in [4], the proposed planner (19) achieves prescribed-time collision-free path planning.
Moreover, the planning algorithm offers an analytical solution without involving any numerical optimization solving, rendering it computationally efficient.
Remark 4.
Although the path planner in (19) is proposed for environments with circular obstacles, it can be easily extended to deal with convex polygonal obstacles, such as rectangle, trapezoid, and triangle. For such obstacles, we can define the augmented obstacle regions $\mathcal{O}_{i}:=\mathcal{O}_{i}^{\ast}\cup\{\mathbf{x}\in\mathbb{R}^{2}\mid{\rm d}_{\mathcal{O}_{i}^{\ast}}(\mathbf{x})<r\}$ and $\mathcal{O}_{i}^{\epsilon}:=\mathcal{O}_{i}^{\ast}\cup\{\mathbf{x}\in\mathbb{R}^{2}\mid{\rm d}_{\mathcal{O}_{i}^{\ast}}(\mathbf{x})<r+\epsilon\}$. The distance function ensures that the boundaries of $\mathcal{O}_{i}$ and $\mathcal{O}_{i}^{\epsilon}$ are smooth, as shown in Fig. 4. In practical implementation, the distance function ${\rm d}_{\mathcal{O}}(\mathbf{x})$ and the bearing vector $\mathbf{b}(\mathbf{x})$ in $\mathbf{h}(\mathbf{x})$ can be obtained locally using on-board LiDARs.
Figure 4: Augmented regions of convex polygonal obstacles.
III-C Prescribed-time tube-following control
To avoid conflicts of notations, the reference trajectory is denoted as $\mathbf{x}_{d}$, which is governed by $\dot{\mathbf{x}}_{d}=\bm{\tau}_{d}$, where $\bm{\tau}_{d}$ is given by (19) but with $\mathbf{x}$ replaced by $\mathbf{x}_{d}$. Define the position tracking error as $\mathbf{x}_{e}:=\mathbf{x}-\mathbf{x}_{d}$. It follows from (3) that
$\dot{\mathbf{x}}_{e}=\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}.$ (21)
To ensure the tracking safety, we impose the following “tube” constraint on $\mathbf{x}_{e}$